Given is a historical time series for job services demand in the prior 6 months.

Month

Demand

1

817

2

736

3

816

4

718

5

805

6

785


Use the table below to answer all questions:

Month

Demand

Forecast

1

817

F1

2

736

F2

3

816

F3

4

718

F4

5

805

F5

6

785

F6

7

F7


If the demands forecast is Not Possible (NP) to generate for a particular month, then you must type in NP as your answer.

a) The F3 by using Naïve forecasting method =  
b) The F3 by using 2 months moving average =  (in 2 decimal places)
c) The F4 by using Naïve forecasting method =  
d) The F4 by using 4 months moving average =  
e) The F6 by using 3 months moving average =  (in 2 decimal places)
f) The F7 by using 4 months moving average =  (in 2 decimal places)
Answer #1

a) The F3 by using Naive forecasting method = Actual Demand of Month 2 = 736

b) The F3 by using 2 months moving average = Average of Actual Demand of Month 1&2 = (817 + 736)/2 = 776.50

c) The F4 by using Naive forecasting method = Actual Demand of Month 3 = 816

d) The F4 by using 4 months moving average = NP

as we need at least the demand of previous 4 months

e) The F6 by using 3 months moving average = Average of Actual demand of previous 3 months (Month 3, 4 & 5) = (816 + 718 + 805)/3 = 779.67

f) The F7 by using 4 months moving average = Average of Actual demand of previous 4 months (Month 3, 4 , 5 & 6) = (816 + 718 + 805 + 785)/4 = 781